# Compact kinks in a modified Christ-Lee model

### v1

@@ I Introduction
Topological solitons appear in several areas of physics, including effective models of particle and nuclear physics, condensed matter systems and cosmological settings [1, 2]. Specially interesting is the case of scalar fields in $1+1$ dimensions, where the presence of multiple vacua is sufficient to allow for the existence of topological solitons called kinks. Topological kinks and their interactions constitute a present-time hectic research area in mathematical physics [3, 4, 5, 6, 7, 8, 9]. Even in the prototypical double well $\phi^{4}$ model, several key features of kink collisions – such as the resonance windows and the value of the critical velocity – have only recently been explained [4, 5, 6]. Similar advances followed in other models [7, 8, 9].
@@
However, those studies have focused on non-analytic potentials given by piece-wise quadratic expressions, with either periodic or double well shapes. In this work, we study topological kinks in models with non-analytic potentials given by piece-wise expressions of higher power orders, including rational and irrational powers, and with different vacuum structures. We propose a fairly general bi-parametrized potential, V-shaped precisely at its minima, possessing a rich vacuum structure similar to the Christ-Lee model, including cases of false vacuum. Besides assuring generality, the two control parameters provide a convenient tool for classification of our analytic and numerical results, easing comparison with previous particular results in the literature. After obtaining the static kink solutions, we discuss its dynamics, analyzing both its internal modes and dynamics in collision processes. In this way, we aim to provide a comprehensive and systematic analysis of compact kinks and their interactions in a broader context, so that general features of compactons can be distinguished from specificities of particular models.
@@ II Model
We shall construct a non-analytic bi-parametrized self-interacting potential inspired by the standard Christ-Lee model [26], which is a scalar field model in $1+1$ dimensions with Lagrangian density
@@
has a parameter $\epsilon$ characterizing its vacuum structure. For $\epsilon=0$, the potential $V_{CL}(\phi)$ gives the prototypical $\phi^{6}$ triple well with three global minima at $\phi\in\{-1,0,1\}$, for $0<\epsilon<1/\sqrt{2}$ the vacuum at $\phi=0$ is false, and for $\epsilon\geq 1/\sqrt{2}$ the potential only has two minima. At the $\epsilon\to\infty$ limit, the Christ-Lee model reduces to the well-known $\phi^{4}$ double well model. As the Christ-Lee potential approaches each minimum parabolically, its solutions are of infinite range, featuring exponential tails. The topological kink solutions and their interactions in the Christ-Lee model have already been studied [27, 28, 7].
@@
in which we impose $\operatorname{sgn}0=0$. This convention agrees with the definition of derivative in the distributional sense and guarantees that the constant field configuration equal to a minimum of the potential corresponds to a solution of the field equation. Since the potential has a V-shape around its minima, the solutions of the field equations have compact support, i.e. are different from the vacuum only inside compact regions of space. From the derivative expression (3) we see that the condition $n\geq 1/2$ is necessary for the one-sided derivatives at $\phi=0$ to be finite.
@@
  • If $10<\epsilon<$ 1, the potential has a local minimum at $\phi=0$, so the model has two true vacua $\phi=\pm 1$ and a false vacuum $\phi=0$.
@@ III Static kinks
for some point $x_{0}$. The solutions for the upper sign have positive $\mathcal{Q}$ and are called kinks and solutions for the lower sign have negative $\mathcal{Q}$ and are called antikinks. Since the field equation has spatial reflection symmetry, for each kink profile $\phi(x)$, the corresponding antikink is given simply by $\phi(-x)$. When there is only two global minima, the model has a single kink and antikink solutions. However, for the triple well cases, we have a kink interpolating the vacua $-1$ and 00 and a second one interpolating the vacua 00 and $1$, together with their corresponding antikinks. We label these kinks and antikinks by the pairs of vacua they interpolate, i.e. $(\phi_{L},\phi_{R})$ where $\phi_{L/R}\in\{-1,0,1\}$.
@@
$\rho(t,x)=\frac{1}{2}\big{(}\partial_{t}\phi(t,x)\big{)}^{2}+\frac{1}{2}\big{(}\partial_{x}\phi(t,x)\big{)}^{2}+V_{n,\epsilon}\big{(}\phi(t,x)\big{)}$ (6)
@@ III.1 $(-1,0)$ and $(0,1)$ kinks
We will first discuss the case of the kinks interpolating the pairs of vacua $(-1,0)$ and $(0,1)$. These kink solutions are possible for the cases $\epsilon=0$, when $\phi=0$ is global minima of the potential, together with $\phi=\pm 1$. The potential has a triple well shape, similar to the $\phi^{6}$ model [31, 32, 33], and is given by the expression
@@
For the case $\epsilon=0$ it is enough to discuss the $(0,1)$ kink, since the $(-1,0)$ can be obtained as $\phi_{(-1,0)}(x)=-\phi_{(0,1)}(-x)$ due to the symmetry of the potential under $\phi\to-\phi$.
For the $(0,1)$ sector, the BPS equation yields the following equation for the kink profile:
@@
The $(0,1)$ kink has support of size
@@
which is a monotonically decreasing function on $n$, with limiting values $L_{(0,1)}(1/2,0)=\pi$ and $L_{(0,1)}(\infty,0)=2$. The $(0,1)$ kink has mass
@@
which increases monotonically with $n$ and has limiting values $M_{(0,1)}(1/2,0)=\pi/8$ and $M_{(0,1)}(\infty,0)=2/3$. The $(-1,0)$ kink will have the same support size and mass.
Note that the $(-1,0)$ and $(0,1)$ kinks are only symmetric under reflection for the special case $n=1/2$. For larger $n$, the potential reaches the vacuum $\phi=0$ differently from the $\phi=\pm 1$ vacua. This is reflected in the kink profiles approach to each vacuum. As $n$ increases, the kink profile approaches $\phi=\pm 1$ in an increasingly sharp way, while the approach to $\phi=0$ has the same behavior for all $n$.
@@ III.2 $(-1,1)$ kinks
For the $\epsilon=0$ case, the compactness of $\phi_{(0,1)}(x)$ allows us to build a larger kink for the $(-1,1)$ sector as a superposition
@@
provided that $a\geq L_{(0,1)}(n,0)/2$, i.e. that the $(-1,0)$ and $(0,1)$ kinks do not overlap. In this case, we refer to the $(-1,0)$ and $(0,1)$ kinks as subkinks. The size of the $(-1,1)$ kink is $L_{(0,1)}(n,0)+2a\geq 2L_{(0,1)}(n,0)$. By convention, we take $a=L_{(0,1)}(n,0)/2$, so that the subkinks have zero separation. For larger separations we treat each subkink as individual kinks. Since the $(-1,0)$ and $(0,1)$ subkinks do not interact when their supports do not overlap, the mass of the $(-1,1)$ kink is simply $2M_{(0,1)}(n,0)$.
For $\epsilon>0$ we do not have $(-1,0)$ and $(0,1)$ kinks, so the $(-1,1)$ kink has to be obtained by directly solving the BPS equation with the appropriate boundary conditions. Note this includes both the cases when $\phi=0$ is a local minimum, or false vacuum, and the cases when $\phi=0$ is a maximum of the potential. In general, the integral form of the BPS equation (4) has no closed form solution, so the kink profile has to be numerically obtained. We present plots for the kink profile in figure 4.
@@
Due to the presence of the false vacuum, the $(-1,1)$ kink for small $\epsilon>0$ can be described as predominantly a superposition of two subkinks in the $(-1,0)$ and $(0,1)$ sectors. This can be clearly seen in the energy density plots in figure 5. For the cases when $\phi=0$ is a local minimum of the potential, the energy density features two maxima, corresponding to the two subkinks. However, as we increase $\epsilon$, these maxima converge until there is only one maximum for $\epsilon=\infty$.
Figure 5: Energy density $\rho(x)$ for the $(-1,1)$ kink.
@@
Figure 6: Support size $L_{(-1,1)}(n,\epsilon)$ and mass $M_{(-1,1)}(n,\epsilon)$ for the $(-1,1)$ kink as functions of $n/(1+n)$ and $\epsilon/(1+\epsilon)$. We change variables for the plot to be able to present the full dependence on the parameters, including the limits towards infinity.
@@ IV Kink internal modes
In this section, we start to analyze the dynamics of our compact kinks. In particular, we look for space-time dependent solutions of the field equation by considering small oscillating perturbations on top of previous static kink profile, so that the field can be written as
@@
where $\chi(x)$ and $\omega$ are the perturbation profile and its angular frequency, respectively. The corresponding solutions for the perturbation profile $\chi(x)$ are known as the kink internal modes. Following the standard procedure, we plug the field expression (9) into the Euler-Lagrange equation (2) and perform a Taylor expansion in the potential derivative. Since the perturbation is small, we keep only the linear terms in $\chi(x)$ in the resulting expression, leading to
@@
for $n>1/2$. In both cases, the stability potential exhibits Dirac deltas with critical points at the vacua $\phi=\pm 1$ for all values of $\epsilon$. Those delta functions lead to infinite barriers in the stability potential for the regions where the kink profile is equal to ±$\pm 1$.
@@ IV.1 Modes of $(0,1)$ kinks
Starting with the $(0,1)$ kinks for the case $\epsilon=0$, we calculated the stability potential numerically and presented its graph in figure 7. For the general case, the stability potential can be obtained only numerically, as we do not have closed form expressions for the kink profile. Moreover, since the kink profile depends on the potential parameter $n$, so does the stability potential. Due to the Dirac distributions appearing in $V^{\prime\prime}(\phi)$, the stability potential displays infinite barriers at the kink borders. This causes the internal modes to also be compact, with the same support as the static kink solution.
Figure 7: Stability potential $U_{s}(x)$ for the $(0,1)$ kinks with different values of $n$.
@@
Figure 8: Frequency $\omega$ of the internal modes of the $(0,1)$ kinks as functions of $n$.
Figure 9: Profile of first three non-zeroth internal modes of the $(0,1)$ kinks for different values of $n$.
@@ IV.2 Modes of $(-1,1)$ kinks
For the $(-1,1)$ kinks, the Dirac deltas in $V^{\prime\prime}(\phi)$ at $\phi=\pm 1$ give rise to infinite walls in the stability potential at the ends of the kink support. Since the kink profile has the value $\phi=0$ at $x=0$, the stability potential can also feature a Dirac delta at that point. In particular, for $n=1/2$, we have a delta barrier for $\epsilon<1$, and a delta well for $\epsilon>1$. For $n=1/2$ and $\epsilon=0$, the potential has a smooth maximum at $\phi=0$, therefore neither barrier nor well show up at $x=0$. Meanwhile, for $n>1/2$ with finite $\epsilon$, the stability potential features a delta barrier at $x=0$. In the limit $\epsilon\rightarrow\infty$, when $n>1/2$, the delta at $\phi=0$ does not contribute. We present plots of the stability potential in figure 10.
Figure 10: Stability potential $U_{s}(x)$ for the $(-1,1)$ kinks with different values of $n$.
We numerically solved the eigenfunction equation (10) to obtain the profiles and the frequencies for the modes of the $(-1,1)$ kinks, which are presented in figures 1113. We observe that for $\epsilon=0$, when the $(-1,1)$ kink is simply a superposition of the $(-1,0)$ and $(0,1)$ subkinks, the internal modes are also superposition of the subkinks internal modes, in either symmetric or antisymmetric fashion, as can be observed in the first row of figure 11. This means that the frequencies’ spectrum will be degenerated, with $\omega_{2k}=\omega_{2k+1}$ for each $k\in\{0,1,2,\ldots\}$. However, as we increase the value of $\epsilon$, the degeneracy is broken, with the frequencies increasing as we increase $\epsilon$, as can be seen for selected values of $n$ in figure 12. Note that the lowest frequency remains at $\omega_{0}=0$, corresponding to the translational mode. We also look at the dependency on the two parameters together in figure 13. The first non-zero mode $\omega_{1}$ has a non-monotonic dependence on $n$, as evidenced by the level curves.
Figure 11: Profile of first four internal modes of the $(-1,1)$ kinks for different values of $n$ and $\epsilon$.
Figure 12: Frequency $\omega$ of the internal modes of the $(-1,1)$ kinks as functions of $\epsilon/(1+\epsilon)$ for some values of $n$.
Figure 13: First three non-zero frequencies $\omega_{k}$ of the modes of $(-1,1)$ kinks as functions of $\epsilon/(1+\epsilon)$ and $n$. The solid lines are level curves.

### v2

@@ I Introduction
Topological solitons appear in several areas of physics, including effective models of particle and nuclear physics, condensed matter systems and cosmological settings [1, 2]. Specially interesting is the case of scalar fields in $+1$ 1 dimensions, where the presence of multiple vacua is sufficient to allow for the existence of topological solitons called kinks. Topological kinks and their interactions constitute a present-time hectic research area in mathematical physics [3, 4, 5, 6, 7, 8, 9]. Even in the prototypical double well $\phi^{4}$ model, several key features of kink collisions – such as the resonance windows and the value of the critical velocity – have only recently been explained [4, 5, 6]. Similar advances followed in other models [7, 8, 9].
@@
However, those studies have focused on non-analytic potentials given by piece-wise quadratic expressions, with either periodic or double well shapes. There has also been some interest in models featuring semi-compact kinks [26]. Meanwhile, in models with smooth potentials, topological kink interactions have been studied for potentials of higher-order polynomials [27, 28, 29, 30, 31], as well as models featuring false vacua [32, 33, 34]. Indeed, the dynamics of non-compact topological kinks has been scrutinized in a wide range of scenarios, distinguishing model particularities from more general features. For example, the spectra of internal modes varies from model to model, strongly affecting the resonance windows in kink-antikink collisions. Similarly, the overall shape of the potential changes the relative size of the core, skin, and tail regions of the kinks, which also affects their interactions [35]. So far, a similarly wide analysis has been missing for the case of compact kinks.
In this work, we study topological kinks in models with non-analytic potentials given by piece-wise expressions of higher power orders, including rational and irrational powers, and with different vacuum structures. We propose a fairly general bi-parametrized potential, V-shaped precisely at its minima, possessing a rich vacuum structure similar to the Christ-Lee model, including cases of false vacuum. Besides assuring generality, the two control parameters provide a convenient tool for classification of our analytic and numerical results, easing comparison with previous particular results in the literature. After obtaining the static kink solutions, we discuss its dynamics, analyzing both its internal modes and dynamics in collision processes. In this way, we aim to provide a comprehensive and systematic analysis of compact kinks and their interactions in a broader context, so that general features of compactons can be distinguished from specificities of particular models.
@@ II Model
We shall construct a non-analytic bi-parametrized self-interacting potential inspired by the standard Christ-Lee model [36], which is a scalar field model in $+1$ 1 dimensions with Lagrangian density
@@
has a parameter $\epsilon$ characterizing its vacuum structure. For $\epsilon=0$, the potential $V_{CL}(\phi)$ gives the prototypical $\phi^{6}$ triple well with three global minima at $\phi\in\{-1,0,1\}$, for $0<\epsilon<1/\sqrt{2}$ the vacuum at $\phi=0$ is false, and for $\epsilon\geq 1/\sqrt{2}$ the potential only has two minima. At the $\epsilon\to\infty$ limit, the Christ-Lee model reduces to the well-known $\phi^{4}$ double well model. As the Christ-Lee potential approaches each minimum parabolically, its solutions are of infinite range, featuring exponential tails.
@@
in which we impose $\operatorname{sgn}0=0$. This convention agrees with the definition of derivative in the distributional sense and guarantees that the constant field configuration equal to a minimum of the potential corresponds to a solution of the field equation. Since the potential has a V-shape around its minima, the solutions of the field equations have compact support, i.e., are different from the vacuum only inside compact regions of space. From the derivative expression (3), we see that the condition $n\geq 1/2$ is necessary for the one-sided derivatives at $\phi=0$ to be finite.
@@
  • If $0<\epsilon<1$, the potential has a local minimum at $\phi=0$, so the model has two true vacua $\phi=\pm 1$ and a false vacuum $\phi=0$.
@@ III Static kinks
for some point $x_{0}$. The solutions for the upper sign have positive $\mathcal{Q}$ and are called kinks and solutions for the lower sign have negative $\mathcal{Q}$ and are called antikinks. Since the field equation has spatial reflection symmetry, for each kink profile $\phi(x)$, the corresponding antikink is given simply by $\phi(-x)$. When there is only two global minima, the model has a single kink and antikink solutions. However, for the triple well cases, we have a kink interpolating the vacua 1-1- 1 and 0 and a second one interpolating the vacua 0 and 111, together with their corresponding antikinks. We label these kinks and antikinks by the pairs of vacua they interpolate, i.e., $(\phi_{L},\phi_{R})$ where $\phi_{L/R}\in\{-1,0,1\}$.
@@
$\rho(t,x)=\frac{1}{2}\big{(}\partial_{t}\phi(t,x)\big{)}^{2}+\frac{1}{2}\big{(}\partial_{x}\phi(t,x)\big{)}^{2}+V_{n,\epsilon}\big{(}\phi(t,x)\big{)},$ (6)
@@ III.1 $(-1,0)$ and $(0,1)$ kinks
We will first discuss the case of the kinks interpolating the pairs of vacua (1,0)(-1,0)( - 1 , 0 ) and (0,1)(0,1)( 0 , 1 ). These kink solutions are possible for the cases $\epsilon=0$, when $\phi=0$ is a global minimum of the potential together with $\phi=\pm 1$. The potential has a triple well shape, similar to the $\phi^{6}$ model [27, 29, 28], and is given by the expression
@@
For the case $\epsilon=0$, it is enough to discuss the (0,1)(0,1)( 0 , 1 ) kink, since the (1,0)(-1,0)( - 1 , 0 ) can be obtained as $\phi_{(-1,0)}(x)=-\phi_{(0,1)}(-x)$ due to the symmetry of the potential under $\phi\to-\phi$.
For the (0,1)(0,1)( 0 , 1 ) sector, the BPS equation yields the following equation for the kink profile:
@@
The (0,1)(0,1)( 0 , 1 ) kink has support of size
@@
which is a monotonically decreasing function on $n$, with limiting values $L_{(0,1)}(1/2,0)=\pi$ and $L_{(0,1)}(\infty,0)=2$. The (0,1)(0,1)( 0 , 1 ) kink has mass
@@
which increases monotonically with $n$ and has limiting values $M_{(0,1)}(1/2,0)=\pi/8$ and $M_{(0,1)}(\infty,0)=2/3$. The (1,0)(-1,0)( - 1 , 0 ) kink will have the same support size and mass.
Note that the (1,0)(-1,0)( - 1 , 0 ) and (0,1)(0,1)( 0 , 1 ) kinks are only symmetric under reflection for the special case $n=1/2$. For larger $n$, the potential reaches the vacuum $\phi=0$ differently from the $\phi=\pm 1$ vacua. This is reflected in the kink profiles approach to each vacuum. As $n$ increases, the kink profile approaches $\phi=\pm 1$ in an increasingly sharp way, while the approach to $\phi=0$ has the same behavior for all $n$.
@@ III.2 $(-1,1)$ kinks
For the $\epsilon=0$ case, the compactness of $\phi_{(0,1)}(x)$ allows us to build a larger kink for the (1,1)(-1,1)( - 1 , 1 ) sector as a superposition
@@
provided that $a\geq L_{(0,1)}(n,0)/2$, i.e., that the (1,0)(-1,0)( - 1 , 0 ) and (0,1)(0,1)( 0 , 1 ) kinks do not overlap. In this case, we refer to the (1,0)(-1,0)( - 1 , 0 ) and (0,1)(0,1)( 0 , 1 ) kinks as subkinks. The size of the (1,1)(-1,1)( - 1 , 1 ) kink is $L_{(0,1)}(n,0)+2a\geq 2L_{(0,1)}(n,0)$. By convention, we take $a=L_{(0,1)}(n,0)/2$, so that the subkinks have zero separation. For larger separations, we treat each subkink as individual kinks. Since the (1,0)(-1,0)( - 1 , 0 ) and (0,1)(0,1)( 0 , 1 ) subkinks do not interact when their supports do not overlap, the mass of the (1,1)(-1,1)( - 1 , 1 ) kink is simply $2M_{(0,1)}(n,0)$.
For $\epsilon>0$ we do not have (1,0)(-1,0)( - 1 , 0 ) and (0,1)(0,1)( 0 , 1 ) kinks, so the (1,1)(-1,1)( - 1 , 1 ) kink has to be obtained by directly solving the BPS equation with the appropriate boundary conditions. Note that this includes both the cases when $\phi=0$ is a local minimum, or false vacuum, and the cases when $\phi=0$ is a maximum of the potential. In general, the integral form of the BPS equation (4) has no closed form solution, so the kink profile has to be numerically obtained. We present plots for the kink profile in figure 4.
@@
Due to the presence of the false vacuum, the (1,1)(-1,1)( - 1 , 1 ) kink for small $\epsilon>0$ can be described as predominantly a superposition of two subkinks in the (1,0)(-1,0)( - 1 , 0 ) and (0,1)(0,1)( 0 , 1 ) sectors. This can be clearly seen in the energy density plots in figure 5. For the cases when $\phi=0$ is a local minimum of the potential, the energy density features two maxima, corresponding to the two subkinks. However, as we increase $\epsilon$, these maxima converge until there is only one maximum for $\epsilon=\infty$.
Figure 5: Energy density $\rho(x)$ for the (1,1)(-1,1)( - 1 , 1 ) kink.
@@
Figure 6: Support size $L_{(-1,1)}(n,\epsilon)$ and mass $M_{(-1,1)}(n,\epsilon)$ for the (1,1)(-1,1)( - 1 , 1 ) kink as functions of $n/(1+n)$ and $\epsilon/(1+\epsilon)$. We change variables for the plot to be able to present the full dependence on the parameters, including the limits towards infinity.
@@ IV Kink internal modes
In this section, we start to analyze the dynamics of our compact kinks. In particular, we look for space-time dependent solutions of the field equation by considering small oscillating perturbations on top of the previous static kink profiles, so that the field can be written as
@@
where $\chi(x)$ and $\omega$ are the perturbation profile and its angular frequency, respectively. The corresponding solutions for the perturbation profile $\chi(x)$ are known as the kink internal modes. Following the same steps as in the standard procedure, we plug the field expression (9) into the Euler-Lagrange equation (2) and perform a Taylor expansion in the potential derivative. Since the perturbation is small, we keep only the linear terms in $\chi(x)$ in the resulting expression, leading to
@@
for $n>1/2$. In both cases, the stability potential exhibits Dirac deltas with critical points at the vacua $\phi=\pm 1$ for all values of $\epsilon$. Those delta functions lead to infinite barriers in the stability potential for the regions where the kink profile is equal to ±1$\pm$ 1± 1.
@@
Another subtle point relates to the presence of radiation. The infinite wall at the stability potential could lead to the impression that no radiation is possible. However, the linearization procedure implemented here does not account for the full spectrum of excitations around the static kink solution. Large enough perturbations are influenced by higher order terms which, in some cases, can lead to the decoupling of radiation from the topological defect. That radiation is composed of compact oscillating packets known as oscillons. Oscillons are solutions for small amplitude configurations in models with non-analytic potentials [17] that dominate the radiation spectrum of models with non-analytic potentials [22]. In fact, the emission of oscillons by excited compact kinks has also been explicitly observed [24].
@@ IV.1 Modes of $(0,1)$ kinks
Starting with the (0,1)(0,1)( 0 , 1 ) kinks for the case $\epsilon=0$, we calculated the stability potential numerically and presented its graph in figure 7. For the general case, the stability potential can be obtained only numerically, as we do not have closed form expressions for the kink profile. Moreover, since the kink profile depends on the potential parameter $n$, so does the stability potential. Due to the Dirac distributions appearing in $V^{\prime\prime}(\phi)$, the stability potential displays infinite barriers at the kink borders. This causes the internal modes to also be compact, with the same support as the static kink solution.
Figure 7: Stability potential $U_{s}(x)$ for the (0,1)(0,1)( 0 , 1 ) kinks with different values of $n$.
@@
Figure 8: Frequency $\omega$ of the internal modes of the (0,1)(0,1)( 0 , 1 ) kinks as functions of $n$.
Figure 9: Profile of first three non-zeroth internal modes of the (0,1)(0,1)( 0 , 1 ) kinks for different values of $n$.
@@ IV.2 Modes of $(-1,1)$ kinks
For the (1,1)(-1,1)( - 1 , 1 ) kinks, the Dirac deltas in $V^{\prime\prime}(\phi)$ at $\phi=\pm 1$ give rise to infinite walls in the stability potential at the ends of the kink support. Since the kink profile has the value $\phi=0$ at $x=0$, the stability potential can also feature a Dirac delta at that point. In particular, for $n=1/2$, we have a delta barrier for $\epsilon<1$, and a delta well for $\epsilon>1$. For $n=1/2$ and $\epsilon=0$, the potential has a smooth maximum at $\phi=0$, therefore neither barrier nor well show up at $x=0$. Meanwhile, for $n>1/2$ with finite $\epsilon$, the stability potential features a delta barrier at $x=0$. In the limit $\epsilon\rightarrow\infty$, when $n>1/2$, the delta at $\phi=0$ does not contribute. We present plots of the stability potential in figure 10.
Figure 10: Stability potential $U_{s}(x)$ for the (1,1)(-1,1)( - 1 , 1 ) kinks with different values of $n$.
We solved the eigenfunction equation (10) numerically to obtain the profiles and the frequencies for the modes of the (1,1)(-1,1)( - 1 , 1 ) kinks, which are presented in figures 1113. We observe that for $\epsilon=0$, when the (1,1)(-1,1)( - 1 , 1 ) kink is simply a superposition of the (1,0)(-1,0)( - 1 , 0 ) and (0,1)(0,1)( 0 , 1 ) subkinks, the internal modes are also superposition of the subkinks internal modes, in either symmetric or antisymmetric fashion, as can be observed in the first row of figure 11. This means that the frequencies’ spectrum is degenerated, with $\omega_{2k}=\omega_{2k+1}$ for each $k\in\{0,1,2,\ldots\}$. However, as we increase the value of $\epsilon$, the degeneracy is broken, with the frequencies increasing as we increase $\epsilon$, as can be seen for selected values of $n$ in figure 12. Note that the lowest frequency remains at $\omega_{0}=0$, corresponding to the translational mode. We also look at the dependency on the two parameters together in figure 13. The first non-zero mode $\omega_{1}$ has a non-monotonic dependence on $n$, as evidenced by the level curves.
Figure 11: Profile of first four internal modes of the (1,1)(-1,1)( - 1 , 1 ) kinks for different values of $n$ and $\epsilon$.
Figure 12: Frequency $\omega$ of the internal modes of the (1,1)(-1,1)( - 1 , 1 ) kinks as functions of $\epsilon/(1+\epsilon)$ for some values of $n$.
Figure 13: First three non-zero frequencies $\omega_{k}$ of the modes of (1,1)(-1,1)( - 1 , 1 ) kinks as functions of $\epsilon/(1+\epsilon)$ and $n$. The solid lines are level curves.